# Bourbaki degree of pairs of projective surfaces

### 31 Mar 2026 v3

@@ 3 Extreme cases of low initial degree
In this section, we observe that a non-zero section $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ induces a sub-foliation by curves of degree $e+1$ of the foliation ${\mathcal{T}}_{\sigma}(1)$. We derive numerical restrictions for this behavior when the initial degree is extremely low $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two. From this main result, we conclude that ${\mathcal{T}}_{\sigma}$ is $\mu$-stable when $\sigma=(f,g)$ is a sequence with $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$, see 27. We will review some of the theory of foliations by curves in $\mathbb{P}^{3}$ (see, for example, [undefc]).
As explored in [undefj, Section 9], the sheaf ${\mathcal{T}}_{\sigma}(1)$ defines a foliation with a corresponding short exact sequence:
$0\rightarrow{\mathcal{T}}_{\sigma}(1)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is the singular scheme of the associated foliation. Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a non-zero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
@@
defining a subscheme $W\subset\mathbb{P}^{3}$, called the singular scheme of the associated foliation by curves. It has codimension at least two and it is also described by ${\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})\simeq\mathcal{O}_{W}$.
The classification of such foliations by curves of low degree provides the following bounds on their singular schemes:
@@
    ([undefk, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or double line of genus 1-1. In either case, $\deg(W)\leq 2$ or $W$ is zero-dimensional.
@@
    (in preparation, V. Cordeiro) If $e=2$, then $\deg(W)\leq 5$ or is zero-dimensional.
@@
one obtains $4\leq c_{2}(G^{\vee})=11-\deg(W)\leq 11$ (see, for example, [undefc, 4.1, 4.2]). Now, if $G^{\vee}$ is stable, Bogomolov’s inequality says $c_{2}(G^{\vee})\geq 7$. Now, assuming that $G^{\vee}$ is not stable, $h^{0}(G^{\vee}(2))\neq 0$ and one may choose a section to form a sequence
@@
where $Y$ is the vanishing locus of this section. Hence, from additivity one gets $0\leq\deg(Y)=c_{2}(\mathcal{I}_{Y})=c_{2}(G^{\vee})-6$. From the first inequality, we conclude $c_{2}(G^{\vee})\geq 6$, and thus $\deg(W)\leq 5$.
@@
To show $(a)$, we dualize the following short exact sequence, obtained above for $e=1$:
$0\rightarrow\mathcal{I}_{B}(2-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
to get a long exact sequence, after simplifying, of the form:
$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-2)\rightarrow$
$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+2)\rightarrow$
$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $\mathcal{U}$ is defined by the short exact sequence $0\to\mathcal{U}\to\mathcal{O}_{\Gamma_{\sigma}}\to\mathcal{O}_{Y}\to 0$ and $Y$ is the one-dimensional component of $\Gamma\sigma$. Moreover, since $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$, we may consider the final piece of the long exact sequence
$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+2)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $W^{\prime}\subset W$ is a pure one-dimensional subscheme. Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$, and from the classification $\deg(W^{\prime})\leq\deg(W)\leq 2$, hence the result follows.
For $(b)$, we proceed analogously, dualizing the sequence
$0\rightarrow\mathcal{I}_{B}(3-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0$
to obtain a long exact sequence, after simplifying, of the form:
$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-3)\rightarrow$
$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+1)\rightarrow$
$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $\mathcal{U}\hookrightarrow\mathcal{O}_{\Gamma_{\sigma}}\rightarrow\mathcal{O}_{Y}$ and $Y$ is the one-dimensional component of $\Gamma_{\sigma}$. Moreover, $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$ and from the final piece of the sequence
$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+1)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $W^{\prime}\subset W$ is a pure one-dimensional subscheme. Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{O}_{\Gamma_{\sigma}},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$. Since $W^{\prime}$ is a subscheme of $W$, we have $\deg(W^{\prime})\leq\deg(W)$. The classification of foliations by curves of degree $2$ implies $\deg(W)\leq 5$, which completes the proof. ∎
@@
[undeft, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree 22 and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
@@
Items $(a)$ and $(b)$ are in the original paper, and to obtain $(c)$, for $a\geq 1$ we can apply the functor $\operatorname{\mathcal{H}\textit{om}}(-,\omega_{\mathbb{P}^{3}})$ to the short exact sequence in $(ii)$, use that $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ and twist by $a+2$ to obtain a short exact sequence
$0\rightarrow\mathcal{O}_{L}(a)\rightarrow\omega_{B}(a+2)\rightarrow\mathcal{O}_{L}\rightarrow 0$
which coincides with the short sequence in $(b)$, hence we obtain the isomorphism in the claim. For the case $a=0$, $B$ must be either a multiplicity two structure as in $(b)$ with $a=0$ or a union of skew lines. In the first case, the proof follows as before. If $B=L_{1}\cup L_{2}$, then $\omega_{B}\simeq\omega_{L_{1}}\oplus\omega_{L_{2}}$ and the claim follows from the fact $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ in each component. ∎
Remark 26.
The results above can be used to restrict the possible values of numerical invariants. For example, for normal pencils of cubics ($d_{f}=d_{g}=2$), $m(\sigma)=7$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, we obtain $\operatorname{Bour}(\sigma)=2$ and we can use the results to show the only possible third Chern class is $c_{3}({\mathcal{T}}_{\sigma})=8$.
Indeed, from [undefu, Theorem 3.8] for the reflexive sheaf ${\mathcal{T}}_{\sigma}(2)$ (unstable of order $r=1$), $c_{3}({\mathcal{T}}_{\sigma})\leq 10$, and using Remark 6 we obtain $c_{3}({\mathcal{T}}_{\sigma})=2(p_{a}(B)+5)=8-2a$, where $p_{a}(B)=-1-a$. From the short exact sequence
$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(4-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
we obtain the Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(4-a))=10-a$ and
$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=10-2a,$
and additivity of Euler characteristic gives a contradiction whenever $a\neq 0$, so $c_{3}({\mathcal{T}}_{\sigma})=8$ if this occurs.
In the next section, we show the above cannot happen either (see 43).
As another application, we are able to show that sequences $\sigma$ with mixed degrees $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$ cannot be unstable.
@@
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials such that $d_{f}=1$, $d_{g}=2$ and $m(\sigma)=3$. If ${\mathcal{T}}_{\sigma}$ is unstable, then the possible discrete invariants are $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=6$.
Proof.
Let us assume that $e=1$, so that $\operatorname{Bour}(\sigma)=2$. From Theorem A Theorem A, $\sigma$ is incompressible. Using Remark 6 we have for $\deg(B)=2$, $p_{a}(B)=-1-a$ yields $c_{3}({\mathcal{T}}_{\sigma})=6-2a\geq 0$, so $a\in\{-1,0,1,2,3\}$, but from
From the short exact sequence
$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(3-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
we get Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(3-a))=8-a$ and
$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=8-2a,$
which can only coincide for $a=0$. ∎
We note that we have no explicit construction for the unstable case above to occur as an example.
@@ 4 Pencils of cubics and degree 6 curves inside quadric surfaces
In this final section, we show some classification results for pencils of cubics and sequences $\sigma=(f,g)$ with $d_{f}=1$, $d_{g}=2$, which correspond to degree $6$ curves inside quadric surfaces.
The results are derived from the previous sections, Section 2.1 and Section 3, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [undefn], [undefw], [undefo], [undefb] and [undefp].
Two major results which are used are the following ones, both assuming that $\kappa$ is an algebraically closed field of characteristic zero.
Theorem 28 ([undefp], Theorem 1.1).
@@ 4.1 Pencils of cubics
When $e=1$, the formula above implies $m(\sigma)\leq 9$, from the inequality $\operatorname{Bour}(\sigma)\geq 0$.
Applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, assuming $e>1$, we get
$c_{1}(\mathcal{E})=c_{1}({\mathcal{T}}_{\sigma}(1))=-d+2=-2$
and the quantities of the Theorem become
$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(4-4+3)\right\rceil=1$
$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil=\left\lceil\frac{1}{3}(4-12+8)\right\rceil=0.$
Moreover, another important invariant is:
$c_{2}(\mathcal{E})$ $=c_{2}({\mathcal{T}}_{\sigma})-d+1$
$=m_{0}-m(\sigma)-d+1$
$=9-m(\sigma).$
Direct application of the result shows the next pair of propositions:
@@
If $\sigma=(f,g)$ is a normal pencil of cubics with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then $m(\sigma)\leq 9$.
@@
If $\sigma=(f,g)$ is a normal pencil of cubics and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then
  • (a)
    if $8\leq m(\sigma)\leq 9$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
  • (b)
    if $0\leq m(\sigma)<7$, then
    $c_{3}\leq m(\sigma)^{2}-17m(\sigma)+72.$
Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(2)$, a sheaf with $c_{1}=0$ and $c_{2}(\mathcal{E})=9-m(\sigma)\geq 0$ from the previous computations, assuming $e>1$. Then, the Theorem says that, for
$t$ $>\sqrt{3c_{2}+1}-2=\sqrt{27-3m(\sigma)+1}-2$
we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+2))\neq 0$.
Direct application of this bound for high values of $m(\sigma)=6,7,8$ gives minimum $t=2,1,1$ satisfying the inequality above, so we obtain
Proposition 32.
Let $\sigma=(f,g)$ be a normal pencil of cubics. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 4$, and we have the following table of sharper bounds for each possibility of $m=m(\sigma)$:
$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
7 $e\leq 3$
8 $e\leq 3$
@@
This follows from intersection theory for the bundle of principal parts (see [undefi], Proposition $7.1$ and Proposition $7.4$). ∎
@@
Since $\mu({\mathcal{T}}_{\sigma})=-2$, we show for $m(\sigma)\leq 6$ that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$. Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 34, 37 and 39. Let us suppose that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. Since $\sigma$ is not free, we can apply the result Theorem B Theorem B, $(a)$, and conclude $\operatorname{Bour}(\sigma)\leq 2$. On the other hand, we have
@@
if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$ established earlier in Theorem B Theorem B, $(b)$. ∎
@@
The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 30 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$. From the formula of the Bourbaki degree, we obtain $\operatorname{Bour}(\sigma)=0$ in this case, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$. On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, then $e=1$, and the equation $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=9$. ∎
@@
If $m(\sigma)=8$, by 32, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$. If $e=2$, we note that $c_{2}({\mathcal{T}}_{\sigma}(2))=0$ and ${\mathcal{T}}_{\sigma}(2)$ is strictly $\mu$-semistable with $c_{1}=0$, thus it follows from [undefb, Lemma 2.0, (a)] that ${\mathcal{T}}_{\sigma}(2)\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}$. If we assume $e\in{1,3}$, we obtain $\operatorname{Bour}(\sigma)=1$ and, using Remark 6 for $B$ being a line (that is, $p_{a}(B)=0$ and $\deg(B)=1$), we obtain:
$e=1$ $\Rightarrow c_{3}=4$
$e=3$ $\Rightarrow c_{3}=0$
On the other hand, 31 gives $c_{3}({\mathcal{T}}_{\sigma})=0$ for $m(\sigma)=8$, so we obtain a contradiction with the case $e=1$. For the case $e=3$, we would have a nearly-free locally free pencil, which cannot happen by 19.
Conversely, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=8$. ∎
@@
Let $\sigma=(f,g)$ be a nearly free pencil of cubics. Then, the only possible discrete invariants are $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $m(\sigma)=7$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 42).
@@
Using Remark 6 for $p_{a}(B)=0$ and $\deg(B)=1$, we obtain
$c_{3}({\mathcal{T}}_{\sigma})$ $=2p_{a}(B)-2+\deg(B)(4+d-2e)$
$=-2+8-2e=2(3-e).$
On the other hand, if $m(\sigma)=7$, then $e=2$ and the formula for $c_{3}$ in Remark 6 yields $c_{3}({\mathcal{T}}_{\sigma})=2$. As we have observed in Example 40 and before, we must have $m(\sigma)\leq 7$ to be able to obtain $\operatorname{Bour}(\sigma)=1$. On the other hand, assuming $m(\sigma)\leq 6$, we obtain
$\operatorname{Bour}(\sigma)=12-m(\sigma)+e(e-4)\geq 6+e(e-4)\geq 2$
hence $\operatorname{Bour}(\sigma)\neq 1$. ∎
Example 42 (Nearly free pencil of Cubics).
@@
For the prime $\mathfrak{p}_{1}$, we may consider the element $u=(2x_{0}(x_{1}-x_{2})-x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{1}}$ and the elementary operation $R_{1}^{\prime}=uR_{1}$, to obtain the matrix
@@
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
@@
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
For the prime $\mathfrak{p}_{3}=(x_{1},x_{2})$, we take $u=(x_{0}^{2}+x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{3}}^{\times}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}+u[x_{2}x_{3}]R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to obtain

### 27 Apr 2026 v4

@@ 3 Extreme cases of low initial degree
As explored in [11, Section 9], a sequence $\sigma=(f,g)$ defines a codimension-one foliation with a corresponding short exact sequence:
$\mathcal{F}_{\sigma}:0\rightarrow{\mathcal{T}}_{\sigma}(1)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is called the singular scheme of $\mathcal{F}_{\sigma}$. Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a nonzero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
@@
defining a subscheme $W\subset\mathbb{P}^{3}$, called the singular scheme of the associated foliation by curves. It has codimension at least two and it is also described by ${\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})\simeq\mathcal{O}_{W}(e-1)$.
We derive numerical restrictions for $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two. A similar strategy was employed to study distributions on $\mathbb{P}^{3}$ with unstable tangent sheaf in [2], and general considerations about the minimal degree of sub-foliations have been considered in [19].
One may localize the diagram above at a minimal prime of the scheme $B$, arguing as in [6, Proposition 3.4]. Over this prime, both sheaves ${\mathcal{T}}_{\sigma}$ and $\mathbb{T}\mathbb{P}^{3}$ split as sums of line bundles, and we obtain a diagram
$\mathcal{O}_{\mathbb{P}^{3},p}(1-e)$$\mathcal{O}_{\mathbb{P}^{3},p}(1-e)$$\mathcal{O}_{\mathbb{P}^{3},p}(-a)\oplus\mathcal{O}_{\mathbb{P}^{3},p}(-b)$$\bigoplus_{i=1}^{3}\mathcal{O}_{\mathbb{P}^{3},p}(l_{i})$$\mathcal{I}_{B}(e-d+1)_{p}$$G_{p}$$\scriptstyle{\nu}$$\scriptstyle{\nu^{\prime}}$$\scriptstyle{L}$
describing $W$ and $B$ via matrix presentations, as the vanishing set of $\nu^{\prime}$ and $\nu$, respectively. From the composition of matrices $L\cdot\nu=\nu^{\prime}$ we conclude minors of $\nu^{\prime}$ are factors of the minors of $\nu$, and thus $B\subset W$ at every associated prime, hence $B\subset W$ schematically. Now, we turn to the classification of schemes $W$ for foliations by curves on $\mathbb{P}^{3}$ of degrees $e=1,2$:
Lemma 26.
Let
$\mathcal{F}:0\to\mathcal{O}_{\mathbb{P}^{3}}(1-e)\to\mathbb{T}\mathbb{P}^{3}\to G\to 0$
be a foliation by curves of degree $e\geq 0$ in $\mathbb{P}^{3}$, with singular scheme $W\subset\mathbb{P}^{3}$. Then:
@@
    ([12, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or a pair of skew lines, in which case $G^{\vee}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}$. In particular, $\deg(W)\leq 2$.
@@
    (in preparation, V. Cordeiro) If $e=2$, then $\deg(W)\leq 5$.
Proof.
@@
one obtains $c_{2}(G^{\vee})=11-\deg(W)\leq 11$ (see, for example, [5, 4.1, 4.2]).If $G^{\vee}$ is stable, Bogomolov’s inequality says $c_{2}(G^{\vee})\geq 7$, hence $\deg(W)\leq 4$. Now, assuming that $G^{\vee}$ is not stable, $h^{0}(G^{\vee}(2))\neq 0$ and one may choose a nonzero section to form a sequence
@@
where $Y$ is the vanishing locus of this section. From the sequence, we get $0\leq\deg(Y)=c_{2}(\mathcal{I}_{Y})=c_{2}(G^{\vee})-6$, so $c_{2}(G^{\vee})\geq 6$, and $\deg(W)\leq 5$. ∎
@@
The result follows from the schematic inclusion $B\subset W$ and from the classification results for the singular scheme $W$, since $\deg(B)\leq\deg(W)$. ∎
@@
[20, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree $2$ and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
@@
For the case $e=1$ and $\operatorname{Bour}(\sigma)=2$, one gets $B=W$ and we can predict the Chern classes for the logarithmic sheaf, by the following proposition. We note that we have not been able to construct examples of sequences of this kind.
Proposition 28.
Let $\sigma=(f,g)$ be a normal sequence with $e=1$ and $\operatorname{Bour}(\sigma)=2$. Then, the Chern classes of ${\mathcal{T}}_{\sigma}$ are (d,d+1,2d)(-d,d+1,2d).
@@
From the formulas $c_{2}({\mathcal{T}}_{\sigma})=m_{0}-m(\sigma)$ and $\operatorname{Bour}(\sigma)=2$, we get $c_{2}({\mathcal{T}}_{\sigma})=d+1$. Since $B=W$ is a pair of skew lines in this case, $p_{a}(B)=-1$ and from
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e)$
it follows that $c_{3}({\mathcal{T}}_{\sigma})=2d$. ∎
The last proposition of the section also uses the foliation structure, and it follows from [6, Proposition B].
@@
Let $\sigma=(f,g)$ be a normal sequence with $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. If $\sigma$ is not free, then ${\mathcal{T}}_{\sigma}$ is not locally free.
@@ 4 Pencils of cubics and degree 6 curves inside quadric surfaces
In this final section, we show some classification results for pencils of cubics and sequences $\sigma=(f,g)$ with $d_{f}=1$, $d_{g}=2$, which correspond to degree $6$ curves inside quadric surfaces. The results are derived from the previous sections, Section 2.1 and Section 3, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [14], [22], [15], [4] and [16]. We recall two major results:
Theorem 30 ([16], Theorem 1.1).
@@ 4.1 Pencils of cubics
When $e=1$, $\operatorname{Bour}(\sigma)\geq 0$ gives $m(\sigma)\leq 9$. Applying Theorem 30 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, assuming $e>1$, we get $c_{1}(\mathcal{E})=-2$ and the quantities of the Theorem become $A=1$ and $B=2$. Since $c_{2}(\mathcal{E})=9-m(\sigma)$ direct application of the result shows the next proposition.
@@
If $\sigma=(f,g)$ is a normal pencil of cubics and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then $m(\sigma)\leq 8$. Furthermore:
  • (a)
    if $7\leq m(\sigma)\leq 8$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
  • (b)
    if $0\leq m(\sigma)<7$, then $c_{3}\leq m(\sigma)^{2}-17m(\sigma)+72$.
Next, we apply Theorem 31 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(2)$, a sheaf with $c_{1}=0$ and $c_{2}(\mathcal{E})=8-m(\sigma)\geq 0$, assuming $e>1$. Then, for $t>\sqrt{3c_{2}+1}-2=\sqrt{24-3m(\sigma)+1}-2$ we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+2))\neq 0$. Direct application of this bound for high values of $m(\sigma)=6,7,8$ gives the following result.
@@
Let $\sigma=(f,g)$ be a normal pencil of cubics. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 4$. If $m(\sigma)=6,7$, $e\leq 3$, and if $m(\sigma)=8$, $e\leq 2$.
@@
This follows from intersection theory for the bundle of principal parts (see [10], Proposition $7.1$ and Proposition $7.4$). ∎
@@
Since $\mu({\mathcal{T}}_{\sigma})=-2$, we show for $m(\sigma)\leq 6$ that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$. Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 35, 38 and 40. Let us suppose that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. Since $\sigma$ is not free, we can apply the result Theorem B Theorem B, $(a)$, and conclude $\operatorname{Bour}(\sigma)\leq 2$. On the other hand, we have
@@
if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$. ∎
@@
The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 32 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$. From the formula of the Bourbaki degree, we obtain $\operatorname{Bour}(\sigma)=0$ in this case, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$. On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, then $e=1$, and the equation $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=9$. ∎
@@
If $m(\sigma)=8$, by $3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 2$, and by 32 it follows that ${\mathcal{T}}_{\sigma}$ is locally free. If $e=1$, then $\operatorname{Bour}(\sigma)=1$, but by 20 nearly-free sequences are never locally free, so this gives a contradiction. Since $\sigma$ is incompressible, it follows that $e=2$, and for $\operatorname{Bour}(\sigma)=0$ we obtain ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}$. Conversely, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ gives $m(\sigma)=8$. ∎
@@
Let $\sigma=(f,g)$ be a nearly-free pencil of cubics. Then, the only possible discrete invariants are $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $m(\sigma)=7$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 43).
@@
By the previous results, $m(\sigma)\leq 7$. But, if $m(\sigma)\leq 6$, we get $\operatorname{Bour}(\sigma)\geq 2$, so $\sigma$ is not nearly free. For $m(\sigma)=7$, any value of $e\neq 2$ gives $\operatorname{Bour}(\sigma)\neq 1$. Thus, it follows that $e=2$. Moreover, $p_{a}(B)=0$ and $\deg(B)=1$ imply $c_{2}({\mathcal{T}}_{\sigma})=3$. ∎
Example 43 (Nearly-free pencil of Cubics).
@@
For the prime $\mathfrak{p}_{1}$, we may consider the element $u=(2x_{0}(x_{1}-x_{2})-x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{1}}$ and an elementary operation to obtain the matrix
@@
so that $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
@@
so that $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
For the prime $\mathfrak{p}_{3}=(x_{1},x_{2})$, we take $u=(x_{0}^{2}+x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{3}}^{\times}$ to perform elementary operations and obtain